### v1

@@ 3.2 Definitions of the Q parameters
The filters GLIMPSE [8], WISE W3, and MSX A contain the 10 m silicate emission. The MSX D and WISE W4 filters appear to be particularly useful as they are not affected by the silicate emission, and show small variations of values in the three curves.
@@
3.3 Q-intrinsic-colours curves for stars with known A(int)
Using the stars by Messineo et al. (2018) as a reference sample, a series of equations is set up to express the intrinsic colours as a function of the parameters.
The parameter Q correlates with the dereddened colour, e.g. (see Fig. 1). The fits are made with third-grade polynomial fits. It follows that
The intrinsic colours are calculated with the median A values of the surrounding stars (Messineo et al. 2005), which are rescaled to a power law with an index of , and the assumed effective extinction ratios. In Messineo et al. (2004) and Messineo et al. (2005), it is shown that the SiO masing stars are rare and the majority of the surrounding stars are normal giants. The large interval spanned by Q is due to the intrinsic colour range of the stars (stellar SEDs). These bright late-type stars are mostly mass-losing Mira AGBs with spectral types later than M4.
For each filter, the constants of the fits are listed in Table 2, when using Curve 3 of Messineo et al. (2005), and in Table 3 for the extinction curve of Gordon et al. (2021).
@@ 3.3 Q-intrinsic-colours curves for stars with known A(int)
3.4 A and (K) colours derived with Q and the Q-intrinsic-colours curves.
The Q-intrinsic-colours curves derived in the above section can be used to obtain estimates of the (K) colours for stars with unknown parameters.
The following formula can be used to estimate the interstellar A as a function of the observed and dereddened colors once the intrinsic colors have been identified with Q:
The term results from Eq. 6 and the coefficients from Table 2 or 3 by entering the Q. The term is the observed colour. The extinction ratios are listed in Table 1.
The Curve 3 of Messineo et al. (2005) was used first. The four independent A obtained with the MSX magnitudes were averaged; 0.05 mag. We repeated the calculation with the WISE dataset and the GLIMPSE dataset. The A values from the W3 and W4 WISE bands give =0.03 mag, while the A values from GLIMPSE [5.8] and [8.0] data yield =0.03 mag. There are 135 stars with good photometry and estimates of A with MSX and WISE datasets; the average difference is 0.04 mag and the is 0.20 mag, as shown in Fig. 2. The 111 A estimates of good photometric quality in common between the GLIMPSE and WISE estimates yield =0.15 mag.
Identical A values are measured (within errors) after repeating the full computation using Gordon’s effective ratios of Table 1. The same A values are obtained as long as the dereddened colours and the are performed consistently with the same interstellar curve.
We measured A by analyzing the correlations between Q and the intrinsic colours. The A values of individual stars are obtained independently of the stellar environment, anywhere in the Galactic Disk.
In Fig. 3, we compare the original estimate of A (via the surrounding stars) and the one obtained with the fits and Q. Using the MSX data, the average A(fit)-A(surrounding)=0.14 mag with a mag and a median of 0.07 mag; using the WISE data the average A(fit)-A(surrounding)=0.04 mag with a mag and a median of 0.00 mag; using the GLIMPSE data A(fit)-A(surrounding)=0.05 mag with a mag and a median of 0.03 mag. This method allows for an accuracy of 0.35 mag, which corresponds to the scatter of the data points in the diagram K versus K of Fig. 1 (the near- and mid-infrared measurements are not acquired simultaneously).
@@ 3.5 Estimates of A as A-A
The K colours of naked giants were calculated by synthetic photometry of theoretical spectra. For this purpose, the publicly available NextGen spectra of Allard et al. (2011) with T=3500, 3200, 2600 K, log(g)=1, and solar metallicity were used333The spectra are distributed by the Virtual Observatory SED analyzer (Bayo et al. 2008). The bt-nextgen_agss2009 (gas only) were retrieved.. The synthetic spectra were generated with the PHOENIX atmospheric code in a spherical symmetry. The synthetic spectra reproduce well the TiO and VO molecules, which dominate the optical spectrum, and water vapour and CO molecules in the infrared spectrum. The molecular list was refined to reproduce the atmospheric lines of brown dwarfs well (Allard et al. 2011; Allard 2014). The infrared colours are obtained by convolving the spectrum with the filter and calculating the mean flux density. The adopted zero points come from Cohen et al. (2003, 2MASS), Egan et al. (2003, MSX), Jarrett et al. (2011, WISE), and Fazio et al. (2004, GLIMPSE). The synthetic infrared colours are listed in Table 4.
The average differences between the A(tot) calculated with the synthetic (K) colours and those obtained with the colours of Koornneef (1983) are +0.05 mag with =0.00 mag (M4) and +0.14 mag with =0.00 mag (M6). While the average differences between the A(tot) calculated with the synthetic (K) colours for a 3500 K (M4) and a 3200 K (M5-6) star are +0.015 mag with =0.006 mag.
The A(int) values were re-determined using the fiducial colour sequences (Q, as described in Sec. 3.
Using the derived A and A, we could estimate the envelope extinction A for 324 targets. A=AA ranges from 0 to 2.5 mag, as shown in Fig. 4. The envelope optical depth at 0.55 m, , can be estimated as A and ranges from 0 to 32.
T | [K] | 3500 | 3200 | 2600 |
Spectral type | M4 | M5-6 | M9 | |
@@
3.6 Empirical circumstellar excess colours.
@@
The mean_1 and sigma_1 values are obtained after having dereddened the observed colours with the A values and the interstellar extinction Curve 3 of Messineo et al. (2005).
Excess ratios | mean_1 | sigma_1 |
(Curve3) | (Curve3) | |
0.68 | 0.43 | |
1.04 | 0.17 | |
1.26 | 0.46 | |
0.34 | 0.78 | |
0.72 | 0.53 | |
3.19 | 36.22 | |
0.05 | 2.97 | |
0.30 | 2.91 | |
1.30 | 0.80 |
The mean_1 and sigma_1 values are obtained after having dereddened the observed colours with the A values and the interstellar extinction Curve 3 of Messineo et al. (2005).
As suggested by the kind referee, the naked colours listed in Table 4 can be useful to determine the colour excess ratios and to add constraints on the circumstellar extinction curve.
@@
where is the interstellar-dereddened colour and is the expected colour for a star of similar parameters (e.g. gravity, metallicity, and temperature) but without an envelope. For the values, the synthetic colours are taken from a solar model at 3200 K (Table 4).
In Fig. 5, the E(Ks-[D]) colour excess is plotted against the E(K-[A]), E(Ks-[C]), and E(Ks-[E]) colour excess. The Figs. 6 and 7 show similar diagrams for the WISE and GLIMPSE filters. There is a tight correlation between the MSX excess colours, as well as between the E(Ks-[W4]) and E(Ks-[W3]) excess colours, and the E(Ks-[8.0]) and E(Ks-[5.8]) excess colours. On the contrary, a large scatter appears in the diagrams showing the E(Ks-[W1]), E(Ks-[W2]), E(Ks-[3.6]), and E(Ks-[4.5]) excess colours in Figs. 6 and 7. Indeed, the stellar spectrum of cold stars short-ward of 4 m is dominated by strong molecular bands (Bedijn 1987). Furthermore, the data used here are from a single-epoch, and near-infrared data and mid-infrared were not taken simultaneously. The radial pulsation expands the atmosphere, and creates an extra layer of molecules (CO, HO, CO, and SiO) above the atmosphere. The region from 2.4 to 4 m is dominated by strong water vapour, which is the main source of opacity. The water vapour absorption changes dramatically with the pulsation phase (Matsuura et al. 2002). This extended layer of molecules was analyzed in detail in the work of Matsuura et al. (2002) and Cami (2002). The larger scatter of data points shown in Fig. 6 and 7, with the filters W1 and W2, and with the filters [3.6] and [4.5], indicates the presence of strong absorption by water.
A tighter correlation of the colour excess E(K-D) with E(K-A), E(K-C), and E(K-E), confirms that the filters A, C, D, and E are dominated by dust emission, as well as W3, W4, [5.8], and [8.0]. As already written by Volk & Kwok (1988) for the LRS spectra, “in the region from 8-23 m the emergent spectrum will directly reflect the shape of the opacity function”.
Using Eq. (8), the excess ratios can also be calculated. The median excess ratios and standard deviations are tabulated in Table 5. The excess ratios in Tables 5 and 6 which were calculated with the MSX bands, WISE W3 and W4, and GLIMPSE [8.0] bands, appear well described by constant values and can provide constraints about the circumstellar extinction curve of the sampled O-rich AGB stars (e.g., Bedijn 1987; Volk & Kwok 1988; Ossenkopf et al. 1992).
@@ 4 A grid of black body models with mass loss
We assume a fixed temperature of 3200 K (M6), the astronomical silicate grains of Ossenkopf et al. (1992), and use the DUSTY code of Ivezic et al. (1999) to generate a set of SED models with optical depth, , from 0 to 40. The model consists of a spherical shell with density decreasing with . The dust condensation radius was set to 1000 K.
In the diagram versus K, the generated model describes the data points of Messineo et al. (2018) well, as shown in Fig. 8. In the diagram versus K, the models display smaller than the data points. Indeed, the W3 filter encompasses the 10 m feature and also the 13 m dust bump; these features are not well reproduced by the silicate of Ossenkopf et al. (1992). Indeed, a much better agreement with the colour is found using a mixture of cold and warm silicates and amorphous alumina (Suh 2021). Fig. 8 shows how changing the results in a a colour-colour track, given a stellar temperature and a dust type. Each point on the track represents a specific tau-point. However, when varying the dust size, the scale changes. This can be seen in Fig. 9, where the dereddened (K is plotted against the envelope. A larger maximum size results in a larger for a given colour. In recent literature, the standard distribution of grain sizes, MNR, is usually adopted with maximum sizes from 0.1 to 1. m (e.g. Olofsson et al. 2022; van Loon et al. 2005; Wang et al. 2021) or fixed sizes of 0.1-0.5 m (e.g. Blommaert et al. 2018; Suh 2021).
Circumstellar effective extinction ratios are highly dependent on the maximum grain size. We estimated that using the DUSTY code, the astronomical silicate grains by Ossenkopf et al. (1992), and an MNR distribution from 0.005 to 0.75 m. For a maximum size of 0.45 m, , and when the maximum size is 0.25 m. Three circumstellar extinction curves with the MNR distribution and maximum sizes of 0.75, 0.45, and 0.25 m are plotted in Fig. 10 to illustrate this. The curves show different peak values of the 9.7 feature (1.4, 0.7, and 0.4, respectively) when normalizing at 2.12 m. Compared to a curve with a maximum grain size of 0.25 m, an extinction curve with a maximum grain size of 0.75 m appears to have a lower 9.7 m peak opacity. The same curve appears steeper in the near-infrared when normaled at 9.7 m, than the curve with a maximum grain size of 0.25 m.
In Table 7, theoretical excess ratios are calculated by convolving the curves with a maximum grain size of 0.75 m and 0.25 m with the filter profiles and a cool spectrum by Allard et al. (2011).
A comparison of the theoretical excess ratios of Table 7 with those observationally measured (Table 5), suggests that an MNR distribution with a maximum grain size of 0.25 m produced excess ratios that were smaller than those observed. An MNR distribution with a maximum grain size of 0.75 m provides a better agreement with the data.
Excess ratios | Temp | mean_1 | mean_2 | mean_3 | mean_4 | mean_4 |
[K] | (max_25) | (max_75) | (Ros_0.4) | (Ros_1.4) | (Ros_0.2) | |
3200 | 0.966 | 0.964 | 0.882 | 0.889 | 0.905 | |
3200 | 0.620 | 0.928 | 0.906 | 0.548 | 0.956 | |
3200 | 0.563 | 0.924 | 0.967 | 0.715 | 1.000 | |
3200 | 1.822 | 0.768 | 0.693 | 1.738 | 0.608 | |
3200 | 2.537 | 1.081 | 0.943 | 2.368 | 0.828 | |
3200 | 1.098 | 1.032 | 1.121 | 1.467 | 1.078 | |
3200 | 0.741 | 0.731 | 0.701 | 0.701 | 0.701 | |
3200 | 0.919 | 0.916 | 0.872 | 0.872 | 0.872 | |
3200 | 0.791 | 0.971 | 1.006 | 0.806 | 1.060 | |
2600 | 0.935 | 0.956 | 0.875 | 0.858 | 0.903 | |
2600 | 0.617 | 0.928 | 0.905 | 0.539 | 0.956 | |
2600 | 0.557 | 0.922 | 0.968 | 0.709 | 1.001 | |
2600 | 1.886 | 0.782 | 0.695 | 1.798 | 0.608 | |
2600 | 2.591 | 1.084 | 0.944 | 2.442 | 0.826 | |
2600 | 1.107 | 1.034 | 1.123 | 1.490 | 1.078 | |
2600 | 0.748 | 0.738 | 0.703 | 0.703 | 0.703 | |
2600 | 0.922 | 0.917 | 0.873 | 0.873 | 0.873 | |
2600 | 0.780 | 0.968 | 1.004 | 0.793 | 1.060 |
@@
Circumstellar excess ratios were calculated with the MSX, WISE, and GLIMPSE filter response curves, with a stellar spectrum of 3200 K (Allard et al. 2011), and extinction curves. The first column (max_25) refers to the extinction curve obtained with DUSTY adopting a maximum size of 0.25 m (see text), which is similar to Rosenthal’s curve normalized at 2.12 m and with a silicate peak at 1.4. The second column (max_75) refers to the extinction curve obtained with DUSTY adopting a maximum size of 0.75 m, which is close to Rosenthal’s curve normalized at 2.12 m and with a silicate peak at 0.4. The spectrum of a naked star of 3200 K is used to obtain the naked star colours.
@@
In order to reproduce the observed colours, small values of the maximum grain sizes (0.1-0.25 m) are frequently reported in the literature. For instance, Volk & Kwok (1988) estimate for their sample of IRAS sources using dust grain sizes smaller than 0.25 m. However, as David & Papoular (1990) points out, observational data (IRAS data) only constrains their opacity curve in the mid-infrared region; for the near-infrared curve, the theoretical model by Draine (1985) is used. Bedijn (1987) uses a power law with an index for the opacity curve below 5 m, yielding .

### v5

@@ 3.2 Definitions of the Qλ parameters
The filters GLIMPSE [8], WISE W3, and MSX A contain the 10 m silicate emission. The MSX D and WISE W4 filters appear to be particularly useful as they are not affected by the silicate emission and show small variations of values in the three curves.
@@
In principle, we can use separately equations (1) and (3) and obtain two independent colour estimates. However, the values span a much smaller range (e.g., from to mag) than the values (e.g., from to mag) and the interpolated colours have large errors. The combined colours span a wider range, for instance, ranges from to mag (see Fig. 1). The combined case is used in the following sections.
3.3 Qλ-intrinsic-colours curves for stars with known A(int)
Using the stars by Messineo et al. (2018) as a reference sample, a series of equations is set up to express the intrinsic colours as a function of the parameters. The parameter Qλ is correlated with the dereddened colour, for instance, (see Fig. 1). The fits are made with third-grade polynomial fits. It follows that:
The intrinsic colours are calculated with the median A values of the surrounding stars (Messineo et al. 2005), which are rescaled to a power law with an index of , and the assumed effective extinction ratios. In Messineo et al. (2004) and Messineo et al. (2005), it is shown that the SiO masing stars are rare and the majority of the surrounding stars are normal giants. The large interval spanned by Qλ is due to the intrinsic colour range of the stars (stellar SEDs). These bright late-type stars are mostly mass-losing Mira AGBs with spectral types later than M4.
For each filter, the constants of the fits are listed in Table 2, when using the revised curve 3 of Messineo et al. (2005) and in Table 3 for the extinction curve of Gordon et al. (2021).
@@ 3.3 Qλ-intrinsic-colours curves for stars with known A(int)
Note: Y-Fit = Coef n1 X-axis + Coef n2 (X-axis)2 + Coef n3 (X-axis)3 +Coef n0.
The colours used for the fit are those dereddened with the A from surrounding stars by Messineo et al. (2005) and Messineo et al. (2018) (rescaled to a power law with an index of ) and effective ratios from curve 3 in Table 1.
@@
3.4 A and (Ks) colours derived with Qλ and the Qλ-intrinsic-colours curves.
The Qλ-intrinsic-colours curves derived in the above section can be used to obtain estimates of the (Ks colours for stars with unknown parameters.
The following formula can be used to estimate the interstellar A as a function of the observed and dereddened colours once the intrinsic colours have been identified with Qλ:
The term results from Eq. 6 and the coefficients from Tables 2 or 3 by entering the Qλ. The term is the observed colour. The extinction ratios are listed in Table 1.
The revised curve 3 of Messineo et al. (2005) was used first. The four independent A obtained with the MSX magnitudes were averaged; 0.05 mag. We repeated the calculation with the WISE dataset and the GLIMPSE dataset. The A values from the W3 and W4 WISE bands give =0.03 mag, while the A values from GLIMPSE [5.8] and [8.0] data yield =0.03 mag. There are 135 stars with good photometry and estimates of A with MSX and WISE datasets; the average difference is 0.04 mag and the is 0.20 mag, as shown in Fig. 2. The 111 A estimates of good photometric quality in common between the GLIMPSE and WISE estimates yield =0.15 mag.
Identical A values are measured (within errors) after repeating the full computation using Gordon’s effective ratios of Table 1. The same A values are obtained as long as the dereddened colours and the are performed consistently with the same interstellar curve. We measured A by analyzing the correlations between Qλ and the intrinsic colours. The A values of individual stars are obtained independently of the stellar environment, anywhere in the Galactic Disk.
In Fig. 3, we compare the original estimate of A (via the surrounding stars) and the one obtained with the fits and Qλ. Using the MSX data, the average A(fit)-A(surrounding)=0.14 mag with a mag and a median of 0.07 mag; using the WISE data the average A(fit)-A(surrounding)=0.04 mag with a mag and a median of 0.00 mag; using the GLIMPSE data A(fit)-A(surrounding)=0.05 mag with a mag and a median of 0.03 mag. The data points displayed in Fig. 3 were fitted with a Huber regression; for the GLIMPSE and the MSX data points, at A=2 mag the fits indicate values 0.2 mag larger than the equity line, suggesting that the new estimated A magnitudes may be 0.2 mag larger than those estimated by Messineo et al. (2005). This systematic error could be corrected a posteriori. However, as shown in the discussion section, for Miras at the Galactic center the new method yields A values that are 0.1 mag smaller than comparison values from the literature. The regression line for the WISE dataset shows a larger offset from the equity line of 0.4 mag at A= 2 mag. This method allows for an accuracy of mag, which corresponds to the scatter of the data points in the diagram Ks versus Ks of Fig. 1 (the NIR and MIR measurements were not acquired simultaneously).
@@ 3.5 Estimates of A as A-A
The Ks colours of naked giants were calculated by synthetic photometry of theoretical spectra. For this purpose, the publicly available NextGen spectra of Allard et al. (2011) with Teff=3500, 3200, 2600 K, log10(g)=1, and solar metallicity were used444The spectra are distributed by the Virtual Observatory SED analyzer (Bayo et al. 2008). The bt-nextgen_agss2009 (gas only) were retrieved.. The synthetic spectra were generated with the PHOENIX atmospheric code considering spherical symmetry. The synthetic spectra reproduce the TiO and VO molecules, which dominate the optical spectrum, and water vapour and CO molecules in the infrared spectrum. The molecular list was refined to reproduce the atmospheric lines of brown dwarfs (Allard et al. 2011; Allard 2014). The infrared colours are obtained by convolving the spectrum with the filter and calculating the mean flux density. The adopted zero points come from Cohen et al. (2003, 2MASS), Egan et al. (2003, MSX), Jarrett et al. (2011, WISE), and Fazio et al. (2004, GLIMPSE). The synthetic infrared colours are listed in Table 4.
The average differences between the A(tot) calculated with the synthetic (Ks) colours and those obtained with the colours of Koornneef (1983) are +0.05 mag with =0.00 mag (M4) and +0.14 mag with =0.00 mag (M6). While the average differences between the A(tot) calculated with the synthetic (Ks) colours for a 3500 K (M4) and a 3200 K (M5-6) star are +0.015 mag with =0.006 mag. The A(int) values were re-determined using the fiducial color sequences Q, as described in Sec. 3.
Using the derived A and A, we could estimate the envelope extinction A for 324 targets. A=AA ranges from 0 to 2.5 mag, as shown in Fig. 4. The envelope’s optical depth at 0.55 m, , ranges from 0 to 32 when estimating it as: A555 AV/A= 12.99 is appropriated for an interstellar infrared power law of index = and RV = 3.1 (Messineo et al. 2005). This conversion factor is close to the 12.82 obtained by Wang & Chen (2019) for Galactic extinction. As shown in the next section, for a circumstellar extinction law, the actual A0.55/A2.2 ratio depends on the adopted grain sizes. It is 15.7 for a maximum size of 0.25 m and 3.7 for a maximum size of 0.75 m..
Teff | [K] | 3500 | 3200 | 2600 |
Spectral typea | M4 | M5-6 | M9c | |
@@
3.6 Empirical circumstellar excess colors.
@@
The mean_1 and sigma_1 values are obtained after having dereddened the observed colours with the A values and the revised Curve 3 of Messineo et al. (2005).
Excess ratios | mean_1 | sigma_1 | mean_2 | sigma_2 |
(Curve3) | (Curve3) | (Gordon) | (Gordon) | |
0.68 | 0.43 | 0.65 | 0.32 | |
1.04 | 0.17 | 0.97 | 0.05 | |
1.26 | 0.46 | 1.22 | 0.24 | |
0.34 | 0.78 | 0.27 | 0.39 | |
0.72 | 0.53 | 0.44 | 0.51 | |
3.19 | 36.22 | 1.54 | 1.77 | |
0.05 | 2.97 | 0.45 | 4.69 | |
0.30 | 2.91 | 0.43 | 3.93 | |
1.30 | 0.80 | 1.18 | 1.46 |
The mean_1 and sigma_1 values are obtained after having dereddened the observed colors with the A values and the revised curve 3 of Messineo et al. (2005).
The naked colours listed in Table 4 can be useful for determining the colour excess ratios and to set constraints on the circumstellar extinction curve.
@@
where is the interstellar-dereddened colour and is the expected color for a star of similar parameters (e.g., gravity, metallicity, and temperature), but without an envelope. For the values, the synthetic colors are taken from a stellar model with a temperature of 3200 K (Table 4).
In Fig. 5, the E(Ks-[D]) colour excess is plotted against the E(K-[A]), E(Ks-[C]), and E(Ks-[E]) color excess. Figs. 6 and 7 show similar diagrams for the WISE and GLIMPSE filters. There is a tight correlation between the MSX excess colours, as well as between the E(Ks-[W4]) and E(Ks-[W3]) excess colors and the E(Ks-[8.0]) and E(Ks-[5.8]) excess colors. On the contrary, a large scatter appears in the diagrams showing the E(Ks-[W1]), E(Ks-[W2]), E(Ks-[3.6]), and E(Ks-[4.5]) excess colours in Figs. 6 and 7. Indeed, the stellar spectrum of cold stars at wavelenghts shorter than 4 m is dominated by strong molecular bands (Bedijn 1987). Furthermore, the data used here are from a single-epoch, and the NIR and MIR data were not taken simultaneously. The radial pulsation expands the atmosphere, and creates an extra layer of molecules (CO, H2O, CO2, and SiO) above the atmosphere. The region from 2.4 to 4 m is dominated by strong water vapour, which is the main source of opacity. The water vapour absorption changes dramatically with the pulsation phase (Matsuura et al. 2002). This extended layer of molecules was analyzed in detail in the work of Matsuura et al. (2002) and Cami (2002). The larger scatter of data points shown in Figs. 6 and 7, with the filters W1 and W2, and with the filters [3.6] and [4.5], indicates the presence of strong absorption by water.
A tighter correlation of the colour excess E(Ks-D) with E(Ks-A), E(Ks-C), and E(Ks-E) confirms that the filters A, C, D, and E are dominated by dust emission, as well as W3, W4, [5.8], and [8.0]. As already reported by Volk & Kwok (1988) for the LRS spectra: “in the region from 8-23 m the emergent spectrum will directly reflect the shape of the opacity function.”
Using Eq. (8), the excess ratios can also be calculated. The median excess ratios and standard deviations are tabulated in Tables 5 and 6. The excess ratios were calculated with the MSX bands, WISE W3 and W4, and GLIMPSE [8.0] bands. They appear to be described by constant values and can provide constraints on the circumstellar extinction curve of the sampled O-rich AGB stars (e.g., Bedijn 1987; Volk & Kwok 1988; Ossenkopf et al. 1992).
@@ 4 A grid of blackbody models with mass loss
Excess ratios | Temp | Ossenkopf et al. (1992)-O-richCold | Suh (1999)-SILC | Rosenthal et al. (2000) | |||||||
curve_1 | curve_2 | curve_3 | curve_4 | curve_5 | curve_6 | curve_7 | curve_8 | curve_9 | |||
[K] | (max_0.25) | (max_0.75) | (max_2.00) | (max_0.25) | (max_0.75) | (max_2.00) | (peak_1.4) | (peak_0.4) | (peak_0.2) | ||
3200 | 0.990 | 0.964 | 0.895 | 1.144 | 0.979 | 0.899 | 0.889 | 0.882 | 0.905 | ||
3200 | 0.624 | 0.905 | 0.854 | 0.966 | 0.979 | 0.923 | 0.548 | 0.906 | 0.956 | ||
3200 | 0.871 | 0.971 | 0.983 | 0.417 | 0.880 | 0.869 | 0.715 | 0.967 | 1.000 | ||
3200 | 1.321 | 0.790 | 0.404 | 1.219 | 0.806 | 0.413 | 1.738 | 0.693 | 0.608 | ||
3200 | 1.674 | 1.046 | 0.677 | 1.499 | 1.040 | 0.671 | 2.368 | 0.943 | 0.828 | ||
3200 | 1.367 | 1.089 | 1.132 | 0.687 | 0.983 | 1.006 | 1.467 | 1.121 | 1.078 | ||
3200 | 0.779 | 0.759 | 0.504 | 0.781 | 0.779 | 0.519 | 0.701 | 0.701 | 0.701 | ||
3200 | 0.921 | 0.923 | 0.733 | 0.898 | 0.925 | 0.734 | 0.872 | 0.872 | 0.872 | ||
3200 | 0.926 | 1.002 | 1.210 | 0.991 | 1.009 | 1.219 | 0.806 | 1.006 | 1.060 | ||
2600 | 0.969 | 0.958 | 0.891 | 1.125 | 0.972 | 0.894 | 0.858 | 0.875 | 0.903 | ||
2600 | 0.614 | 0.905 | 0.853 | 0.969 | 0.979 | 0.923 | 0.539 | 0.905 | 0.956 | ||
2600 | 0.869 | 0.971 | 0.982 | 0.400 | 0.878 | 0.869 | 0.709 | 0.968 | 1.001 | ||
2600 | 1.365 | 0.799 | 0.408 | 1.257 | 0.815 | 0.416 | 1.798 | 0.695 | 0.608 | ||
2600 | 1.726 | 1.051 | 0.680 | 1.539 | 1.043 | 0.672 | 2.442 | 0.944 | 0.826 | ||
2600 | 1.398 | 1.093 | 1.138 | 0.685 | 0.986 | 1.009 | 1.490 | 1.123 | 1.078 | ||
2600 | 0.781 | 0.763 | 0.507 | 0.781 | 0.782 | 0.521 | 0.703 | 0.703 | 0.703 | ||
2600 | 0.920 | 0.924 | 0.734 | 0.898 | 0.926 | 0.736 | 0.873 | 0.873 | 0.873 | ||
2600 | 0.920 | 1.001 | 1.212 | 0.985 | 1.008 | 1.223 | 0.793 | 1.004 | 1.060 |
@@
Circumstellar excess ratios were calculated with the MSX, WISE, and GLIMPSE filter response curves, with a stellar spectrum (Allard et al. 2011) and an extinction curve. The columns curve_1, curve_2, and curve_3 refer to the circumstellar extinction curves obtained with DUSTY and the cold O-rich silicates by Ossenkopf et al. (1992) adopting a maximum size of 0.25, 0.75, and 2.0 m, respectively (see text). The columns curve_4, curve_5, and curve_6 refer to the circumstellar extinction curves obtained with DUSTY and the cold silicates (SILC) by Suh (1999) adopting a maximum size of 0.25, 0.75, and 2.0 m, respectively, (see text). The spectrum of a naked star of 3200 K is used to obtain the naked star colours in the upper part of the table (first 9 rows). In the lower part of the table, the spectrum of a naked star of 2600 K is used. For comparison, the columns curve_7, curve_8, and curve_9 refer to the parametric formula for Galactic interstellar extinction by Rosenthal et al. (2000) normalized at 2.12 m with the silicate peak set at 1.4, 0.4, and 0.2, respectively.
We assumed a fixed temperature of 3200 K (M6), the astronomical silicate grains of Ossenkopf et al. (1992) as well as the silicate grains of Suh (1999). We used the DUSTY code of Ivezic et al. (1999) to generate a set of SED models with optical depth, , from 0 to 40. The model consists of a spherical shell with density decreasing with . The dust condensation radius was set to 1000 K.
In the diagram depicting versus Ks, the models aptly describe the data points of Messineo et al. (2018), as shown in Fig. 8. In the diagram versus Ks, the models with the Ossenkopf’s silicate grains display smaller than the data points, while the models with the cold silicates by Suh (1999) reproduce the observed colours (Suh 2021). Indeed, the W3 filter encompasses the 10 m, as well as the 13 m, and the 18 m dust features. SED models with Suh’s silicates and models with Ossenkopf’s silicates agree short-ward of 10 m; but they disagree around 18 m (Suh 1999). Suh’s warm silicates generate a smaller emission feature at 18 m than the cold silicates (Suh 1999). Suh’s cold silicates generate a stronger absorption feature at 18 m and a stronger continuum longward of 18 m than Ossenkopf’s silicates (Suh 1999) and generate larger W3-W4 colours.
Figure 8 shows how changing the results in a colour-colour track, given a stellar temperature and a dust type. Each point on the track represents a specific tau-point. However, when varying the dust size, the scale changes. This can be seen in Fig. 9, where the dereddened (Ks is plotted against the envelope. A larger maximum size results in a larger for a given colour. In recent literature, the standard distribution of grain sizes, MNR, is usually adopted with maximum sizes from 0.1 to 1. m (e.g., Olofsson et al. 2022; van Loon et al. 2005; Wang et al. 2021) or fixed sizes of 0.1-0.5 m (e.g., Blommaert et al. 2018; Suh 2021).
Circumstellar effective extinction ratios are highly dependent on the maximum grain size. We estimated that using the DUSTY code, the astronomical silicate grains by Ossenkopf et al. (1992), and an MNR distribution from 0.005 to 0.75 m. For a maximum size of 0.45 m, , and when the maximum size is 0.25 m. Four circumstellar extinction curves with the MNR distribution and maximum sizes of 2.00, 0.75, 0.45, and 0.25 m are plotted in Fig. 10 to illustrate this. The curves show different peak values of the 9.7 feature (0.49-0.50, 0.44-0.46, 0.77-0.78, and 1.59-1.61, respectively) when normalizing at 2.12 m. Compared to a curve with a maximum grain size of 0.25 m, an extinction curve with a maximum grain size of 0.75 m appears to have a lower 9.7 m peak opacity. In the near-infrared, the same curve appears located above the curve with a maximum grain size of 0.25 m, when normalized at 9.7 m.
In Table 7, theoretical excess ratios are calculated by convolving the curves with a maximum grain size of 2.00 m, 0.75 m, and 0.25 m with the filter profiles and a cool spectrum by Allard et al. (2011). A comparison of the theoretical excess ratios of Table 7 with those observationally measured (Table 5), suggests that an MNR distribution with a maximum grain size of 0.25 m produces excess ratios that are incompatible with those observed ( much lower than unity and above unity). An MNR distribution with a maximum grain size of 0.75 m provides a better agreement with the data for the MSX data. An MNR distribution with a maximum grain size of 2.00 m also reproduces the trends seen in Table 5 for the WISE and GLIMPSE filters short-ward of 8 m (these filters are affected by molecular absorption).
@@
In order to reproduce the observed colours, small values of the maximum grain sizes (0.1-0.25 m) are frequently reported in the literature. For instance, Volk & Kwok (1988) estimated for their sample of IRAS sources using dust grain sizes smaller than 0.25 m. However, as David & Papoular (1990) points out, observational data (IRAS data) only constrains their opacity curve in the mid-infrared region; for the near-infrared curve, the theoretical model by Draine (1985) is used. Bedijn (1987) uses a power law with an index for the opacity curve below 5 m, yielding .